OpenStudy (anonymous):

Find the value of y for which the line A & B has the given slope m: A(-2, 3), B(4,y), m=-2/3. Thanks

OpenStudy (amistre64):

use the slope intercept form to define the line with the slope and A, then plug in x=4 to solve for the missing y

OpenStudy (amistre64):

\[y=m(x-A_x)+A_y\] \[B_y=m(4-A_x)+A_y\]

OpenStudy (anonymous):

Okay thanks .

OpenStudy (anonymous):

I got -1

OpenStudy (amistre64):

another method is just to use the slope formula A(-2, 3) - B( 4 ,y ) ---------- -6, 3-y \[\frac{y-3}{6}=-\frac23\]

OpenStudy (amistre64):

y-3=-4 when y=-1

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

youre welcome